# Empirical analysis of Binding Precedent efficiency in the Brazilian Supreme Court via Similar Case Retrieval

### v2

@@ 2.2 NLP for legal documents
Moreover, the non-existence of a standardized citation system for the STF allows for what the authors call indirect citations, that is, references to documents that cannot be captured through parsing techniques based on regex. In practice, this means that the interpretation of the authority of a precedent as the number of citations to it at a certain interval of time is insufficient to capture their impact on the legal system.
Finally, the situation is aggravated by the defensive posture assumed by the court to cope with its excessively high demands, with many of the most cited precedents being only procedural decisions of the STF denying to arbitrate on the merit of cases and directing them to other courts. This distinction between decisions in which the precedent use has a significant juridical value in the judge’s juridical reasoning, which we shall call citations by merit, versus those in which the citation is only a procedural justification for dictating the next formal steps of the case, which we call procedural citations, impose a noteworthy constraint in the interpretation of the number of citations of a document as metric for its authority, for a particularly juridically insignificant decision may be frequently cited due to some incidental procedural extract. These constraints are particularly challenging to deal with in the study of the STF: although the vast majority of court decisions cite precedent, most of them lie in the procedural category [34], limiting the ordinary interpretation of authority scores.
The infant precedent citation culture in Brazil and the predominance of procedural-only citation discourage the use of graph techniques in the study of BPs. Additionally, the very use of these precedents as a tool to decrease the frequency of repetitive cases arriving at the STF indicates that degree-based scores are insufficient to describe the effectiveness of the BPs as defense devices, suggesting the use of embedding-based methods.
3 Datasets and methods
@@ 3.1 Datasets
Cases decided at the STF are consolidated in a document, containing the decision’s text, as well as the publication date, the Justice that conducted the process (Justice rapporteur), and the document type (e.g., Complaint or Extraordinary Appeal). The research project Supremo em Números gathered, in text format, more than 2,500,000 of these documents [23]. In this collection, we shall consider the subset of documents that cite a binding precedent, among the 58 precedents edited by STF until 2018. This amounts to 29,743 documents, which we refer to as Dataset #1 (see Table 1). In addition, we shall consider the so-called Dataset #2, consisting of those documents, in the whole collection, labeled as “administrative law”, “criminal law” or “criminal procedure law”, gathering 615,262 documents. These are the categories of the five particular BPs we will study (already presented in Section 2.1). We draw the reader’s attention to the fact that Dataset #1 is not a subset of Dataset #2, though the latter contains approximately 36% of the former.
Dataset #1 Number of docs.
BP 11 830
BP 14 1601
BP 17 856
BP 26 954
BP 37 3984
Other BPs 21,518
Total 29,743
Dataset #2 Number of docs.
Administrative law 463,497
@@
Labels
@@
In the context of Similar Case Retrieval, Dataset #1 is particularly convenient. Indeed, since the 58 binding precedents cover different themes, one decides whether a document deals with a specific theme by reading the corresponding BP it cites. In opposition, no information is available in Dataset #2 regarding the juridical content of the cases. For this reason, we will employ the former to train our models. In the rest of this section, only Dataset #1 will be considered, while the application of our models to Dataset #2 is the content of Section 4.
3.2 Models for Similar Case Retrieval
@@
In this section, we delve into the methodologies employed for analyzing our dataset, focusing on four distinct approaches: TF-IDF, LSTM, BERT, and regex. Together, they faithfully represent some of the most common NLP methods.
@@
Term Frequency-Inverse Document Frequency (TF-IDF) is a vector embedding technique divided into two parts. In the “term-frequency” part, each document $d$ is embedded in the space $\mathbb{R}^{n}$, where $n$ is the number of different words on the corpora, and where the $i$-th component of the vector is the frequency that the $i$-th word of the corpora appears in the document. We denote it by $\text{TF}(i,d)$. The “inverse document frequency” part is another $\mathbb{R}^{n}$ vector, with the $i$-th coordinate given by the logarithm
@@
By itself, TF-IDF is only an embedding, so it must be coupled with some classification algorithm for us to perform Similar Case Retrieval. In this article, we will use three kinds of classification models for the embedded documents: linear Support Vector Machine (SVM), logistic regression, and random forests. We will not explicitly describe these classifiers here, but refer the interested reader to [27].
@@
A recurrent network architecture was also implemented for this task, given the well-known capacity of these models to work with textual data, especially for longer texts [27, 50]. The implemented neural networks were composed of three layers: one embedding layer, followed by two layers consisting of 128 LSTM (Long Short-Term Memory) neurons each, all using relu activation functions. For the embedding layer, tensorflow was used, where the vocabulary (i.e., set of unique words in the corpora) was restricted to the 10,000 most common terms, using 1,000 out-of-bucket embedding for the rest of them. Moreover, because the architecture expects a constant input size of documents, only the first 512 tokens of each document were considered, using additional masking for shorter documents. Although naive, this approach is common when dealing with long text analysis in neural networks. Finally, the output consisted of a single neuron followed by a sigmoid non-linearity. We stress that no pre-training of the network’s weights was used.
BERT
Bidirectional Encoder Representations from Transformer (BERT) [21] is a common Transformer-based model [66] used for several NLP tasks that can benefit from a deep embedding of a word or a sequence, e.g., a text classification. In this model, each token (word or part of a word) is vectorized and their vectors are sequentially updated to incorporate their neighbors’ information. BERT models are released with their weights already optimized in a large amount of text in a self-supervised manner. The output of the model is a vector for each token and one for the entire sentence (corresponding to the first CLS token). These vectors can be used as features for any classification model, e.g., a logistic regression. One can also extend this neural network by adding a downstream neural network layer, like a linear classification layer, and then both the new layer and the rest of the model are trained together in the downstream task. Just like the LSTM implementations, BERT models have a text input size limited to 512 tokens. Because of that, texts were truncated to their first 510 tokens (to include other BERT special tokens).
In our work, we use a Portuguese BERT (BERTimbau BASE [^4] Available at https://huggingface.co/neuralmind/bert-base-portuguese-cased.) model fine-tuned in approximately 16 GB of Brazilian court decision raw data for eight epochs. This data gather decisions from the five most important Brazilian courts: Supremo Tribunal Federal, Superior Tribunal de Justiça, Tribunal de Justiça de São Paulo, Tribunal de Justiça do Rio de Janeiro, and Tribunal Superior Eleitoral.
@@
Regular expression (regex) stands out as a commonly used tool among professionals for retrieving similar cases. In particular, it constitutes the search engine of the official STF website. We designed a regex search for each of the five BPs by selecting, with the help of legal expertise, the most important words in their statements (please refer to Sections 4.2 to 4.6 for the statement and juridical context of each BPs). We draw the reader’s attention to the fact that, as our models are intended to detect, in a subsequent section, the documents before creation of the BPs, we do not use the words Súmula Vinculante or any specific mention of a law. More precisely, we consider:
@@ 3.3 Validation
Our first analysis focuses on Dataset #1, a collection of documents annotated by the binding precedent they cite. We are, therefore, dealing with a binary classification task. As we will see in this section, and as it is expected from the NLP literature, all the models perform well on this task. However, we remind the reader that the main objective of this article lies in the analysis of Dataset #2 (in Section 4), which is unannotated, and where the “scores” will be quite different.
@@
Some preprocessing was applied to the documents before training. For both the TF-IDF and LSTM models, this encompassed the removal of Portuguese-specific accents, and of usual stop words (e.g., a, o, um, mas). Because BERT does not expect this sort of preprocessing, raw texts were given.
Additionally, we want to avoid the models from learning textual cues that mark direct citations to BPs, such as the very term Súmula Vinculante or similar. Therefore, regular expressions were used to substitute these terms by empty strings. Additionally, as our interest consists in detecting possible uses of BPs before their publications, dates were also removed. Once more, BERT models receive a slightly different masking process than the others: each BP citation is identified using regular expressions and is replaced by [MASK] tokens, keeping the same length as the original text.
@@
$\mathrm{Recall}$ $=\mathrm{TP}/\mathrm{P}$
@@
Intuitively, accuracy measures the ratio of correct predictions, while precision measures the accuracy of pointing to class 1, and recall measures how much of class 1 is being correctly predicted. The $F_{1}$ score is the harmonic mean of recall and precision.
Most models return a continuous output, e.g., a probability. In this case, defining when the model is predicting classes 0 or 1 depends on the classification threshold. The Area Under the Precision Recall Curve (AUPRC) swaps a classification threshold extreme values, measuring precision and recall for each threshold, and creates a precision vs. recall curve. The area under the precision recall curve establishes a common performance metric. The results are gathered in LABEL:tab:similar_case_matching.
Table 2: For each BP and each model, we indicate the scores $F_{1}$, precision, recall, and AUPRC (percentage). The best score is shown in bold.
@@
11 TF-IDF + SVM 93.9 98.6 89.6 88.6
TF-IDF + logistic 92.3 100 85.7 86.1
TF-IDF + forest 85.1 100 74.0 74.7
LSTM 87.8 98.4 79.2 78.5
BERT 90.5 94.4 87.0 82.5
regex 91.7 98.5 85.7 84.8
14 TF-IDF + SVM 96.3 97.0 95.6 92.9
TF-IDF + logistic 97.0 98.5 95.6 94.3
TF-IDF + forest 94.1 100 88.9 89.4
LSTM 95.5 96.9 94.1 91.5
BERT 92.1 93.2 91.1 85.3
regex 73.1 97.5 58.5 59.0
17 TF-IDF + SVM 96.7 98.7 94.9 93.7
TF-IDF + logistic 98.0 100 96.2 96.3
TF-IDF + forest 90.1 100 82.1 82.5
LSTM 93.9 100 88.5 88.8
BERT 90.5 95.7 85.9 82.6
regex 88.1 86.4 89.7 77.8
26 TF-IDF + SVM 98.6 100 97.3 97.4
TF-IDF + logistic 98.6 100 97.3 97.4
TF-IDF + forest 95.8 100 91.9 92.1
LSTM 93.7 97.1 90.5 88.2
BERT 94.4 97.1 91.9 89.5
regex 86.2 100 75.7 76.3
37 TF-IDF + SVM 95.2 97.4 93.1 91.4
TF-IDF + logistic 94.5 97.0 92.1 90.3
TF-IDF + forest 85.4 100 74.5 77.4
LSTM 91.7 95.3 88.4 85.5
BERT 90.6 93.6 87.7 83.5
regex 83.4 93.7 75.2 73.2

### v3

@@ 2.2 NLP for legal documents
Moreover, this non-existence of a standardized citation system for the STF allows for what the authors call indirect citations, that is, references to documents that cannot be captured through parsing techniques based on regex. In practice, this means that the interpretation of the authority of a precedent as the number of citations to it at a certain interval of time is insufficient to capture their impact on the legal system.
Finally, the situation is aggravated by the defensive posture assumed by the court to cope with its excessively high demands, with many of the most cited precedents being only procedural decisions of the STF denying to arbitrate on the merit of cases and directing them to other courts. This distinction between decisions where the precedent use has a significant juridical value in the judge’s juridical reasoning, which we shall call citations by merit, versus those in which the citation is only a procedural justification for dictating the next formal steps of the case, which we call procedural citations, impose a noteworthy constraint in the interpretation of the number of citations of a document as metric for its authority, for a particularly juridically insignificant decision may be frequently cited due to some incidental procedural extract. These constraints are specially challenging to deal with in the study of the STF: although the vast majority of court decisions cite precedent, most of them lie in the procedural category [35], limiting the ordinary interpretation of authority scores.
One is then led to conclude that the infant precedent citation culture in Brazil and the predominance of procedural-only citations discourage the use of graph techniques in the study of BPs. Additionally, the very use of these precedents as a tool to decrease the frequency of repetitive cases arriving at the STF indicates that degree-based scores are insufficient to describe the effectiveness of the BPs as defense devices, suggesting the use of embedding-based methods.
3 Training and testing on Dataset #1
@@ 3.1 Datasets
Cases decided at the STF are consolidated in a document, containing the decision’s text, as well as various metadata. The research project Supremo em Números gathered, in text format, more than 2,500,000 of these documents, published between 1989 and 2018 [24]. They were obtained by scraping the official STF website [^3] STF search engine https://portal.stf.jus.br/, where all the documents are publicly accessible.
It is worth noting that the documents follow a typical structure:
  1. Cabeçalho (Header)
    Contains a number of metadata, such as the case identification number, the involved parties, the relator (Justice in charge), the origin of the case (if appealed from a lower court), and its type (classe processual, e.g., Rcl).
  2. Ementa (Abstract)
    A summary of the key points of the decision, highlighting the legal thesis, the constitutional provision at issue, and the most relevant reasoning.
  3. Relatório (Report)
    Description of the procedural history of the case. It summarizes the main facts, arguments, lower court decisions (if any), and other relevant filings.
  4. Fundamentação (Reasoning)
    Detailed legal and argumentative analysis. It presents the interpretation of the relevant legal and constitutional norms, and how precedents apply.
  5. Dispositivo (Ruling)
    The final part of the decision, stating the outcome in clear terms. In Plenary or Panel decisions, the ruling reflects the voting result.
Although one could analyze documents by distinguishing the different parts that compose them, in this article, we have chosen to treat them as a whole. In particular, and as described in Section 3.2, two embedding methods will be considered: bag-of-words models (with TF-IDF) and more advanced NLP methods (LSTM and Longformer).
In addition, each document of the collection is associated with a category (ramo do direito, “branch of law”), such as “Administrative Law” or “Criminal Law”. Although there is no official list published by the STF, these categories align with the well-known branches of Brazilian law. They are used internally, on the official website, to classify laws and documents [^4] One can search for a binding precedent based on its branch at https://portal.stf.jus.br/jurisprudencia/aplicacaosumula.asp, or, similarly, for a document at https://jurisprudencia.stf.jus.br/pages/search.
We draw the reader’s attention to the fact that the metadata presented so far, associated with each document — e.g., the relator, document type, or its branch — is non-ambiguous and does not depend on annotation by an expert. However, we will need to extract other types of information from the documents. This includes, for instance, the state of provenance of the document, the outcome of the decision, the use of a specific law, or more generally the presence of certain words. This will be achieved by a simple regular expression search in the documents. Furthermore, in order to verify the content of certain documents (whether to check that the models have correctly detected a document or to better understand its content), we will have to read them. In this case, the information is potentially ambiguous and we will make this explicit each time.
Construction of our datasets
From the collection of documents gathered by Supremo em Números, we have devised two datasets that will allow us to perform our analysis of binding precedent efficiency. For the first, we have collected, through regular expression searches, all documents that cite a binding precedent, among the 58 precedents edited by the STF until 2018. Then, we have chosen to consider only the ten most cited precedents and to discard the others. The final dataset amounts to 29,743 documents, citing a precedent among BPs 3, 4, 10, 11, 14, 17, 20, 26, 33, and 37. We will refer to it as Dataset #1.
We remind the reader that this article focuses specifically on the precedents 11, 14, 17, 26, and 37. As already presented in Section 2.1, they were chosen for the diversity of the subjects they cover, and the high number of cases they produce. The number of documents corresponding to each of these BPs is given in Table 1. We draw the reader’s attention to the fact that some documents in our collection cite several BPs. More precisely, out of the 29,743 documents, 1116 cite exactly two BPs, and 93 cite three or more. As a consequence, in the table for Dataset #1, the total number of documents is not equal to the sum of the values in the rows above.
In order to take a more comprehensive look at the issues covered by these precedents, we shall also consider the so-called Dataset #2, consisting of those documents, in the whole collection, whose branch is “Administrative Law”, “Criminal Law” or “Criminal Procedure Law”. These are the categories of the five specific BPs under study. In particular, Dataset #1 is a subset of Dataset #2. The number of documents contained in Dataset #2 is 634,068, and we display the number per branch in Table 1.
To provide a last insight into these datasets, Fig. 3 shows the time series of document publication dates. We can see that, in both cases, the trend is increasing, reflecting the congestion of the Supreme Court over time.
Dataset #1
Number of
documents
Of which cite
another BP
BP 11 830 42
BP 14 1601 62
BP 17 856 16
BP 26 954 188
BP 37 3984 895
Other BPs 22,650 1138
Total 29,743 1209
Dataset #2
Number of
documents
Administrative Law 480,903
Criminal Procedure Law 106,537
@@ 5. Dispositivo (Ruling)
Time series of the publication date of documents in Dataset #1 Refer to caption Time series of the publication date of documents in Dataset #2 Refer to caption
Figure 3: Histograms of the number of cases in Dataset #1 (top) and Dataset #2 (bottom). The bins have a length of one year.
Case Classification and labels
In this article, we seek to analyze the influence of binding precedents on Brazilian case law. To do so, our methodology consists of detecting all documents that relate to the theme of the studied BP and analyzing them. These documents may be those that cite the BP, but also those that could use it without citing it explicitly (potential citations), or those that were published before the BP, but could have cited it if the BP already existed. In other words, we are dealing with a Similar Document Retrieval problem, where the cases are those that fall within the scope of the BP. We therefore need to train models to detect potential citations of the precedent.
@@
In this context, Dataset #1 is particularly convenient. Indeed, since the various binding precedents cover different themes, one decides whether a document deals with a specific theme by reading the corresponding BP(s) it cites. In opposition, in Dataset #2, apart from the branch to which the document belongs, no information is available regarding its juridical content. For this reason, we will employ the former to train and test our models. This way, the problem is reduced to that of Case Classification, in the form of a mere binary classification task. More specifically, since we want to train the models not to detect citations but potential applications of the BP, we will mask explicit mentions of the precedent in the training (and testing) data. This approach simulates a situation where the BP does not yet exist. In the rest of this section, only Dataset #1 will be considered.
However, restricting our study to Dataset #1 will not solve our problem, since we would then only obtain documents that explicitly cite some BP. To carry out our program, we, therefore, apply these models to Dataset #2, which contains the document categories in which we expect to detect potential cases of application of the precedents. The application of our models to Dataset #2 will be studied from a Machine Learning perspective in Section 4 and a juridical perspective in Section 5. More precisely, in the former section, we will examine the generalization problem we face when passing from Dataset #1 to Dataset #2, and in the latter, we will use our findings to draw juridical conclusions regarding the BPs considered.
Length of the documents
When it comes to implementing automatic document analysis, an important quantity is the length of the documents. While some methods accept documents with an arbitrary length — for instance, bag-of-words vectorizations —, others accept only a limited input. As an example, the popular transformer-based model BERT treats, in its native form, documents of 512 tokens at most [22]. As we will describe more thoroughly in Section 3.2, we have decided to consider three distinct models: TF-IDF, LSTM, and Longformer. The first embedding accepts documents of any length, but the last two have a limitation, which must be chosen a priori. We have chosen, for both models, a maximal length of 4096 words or tokens.
To get an idea of the impact of this limitation, Fig. 4(a) shows the distribution of document lengths in Dataset #1. In particular, the median length is 1019 words, and very few documents are longer than 4096 words (only $3\%$). That said, the models considered expect a specific preprocessing of the data. For TF-IDF and LSTM, this encompasses the lemmatization of the texts and the removal of Portuguese-specific stop words. For Longformer, a specific tokenizer is used. The distribution of the length of the documents, after the application of these preprocessings, is shown in Figs. 5(a) and 6(a), respectively. Clearly, after lemmatization and stop words removal, the length of the documents is reduced. In particular, the choice of a maximum of 4096 words for LSTM means that almost all documents can be processed in their entirety (only 0.8% exceed the limit). On the other hand, after Longformer’s tokenization, the documents increase in length. Indeed, long or rare words are typically broken into 2 or 3 tokens. Fortunately, and as visualized in Fig. 6(a), only $6\%$ of the tokenized documents have a length greater than 4096. We conclude that our choice of maximum length should not significantly impair the analysis.
Distribution of the number of words in raw documents (no preprocessing) Refer to caption
(a) Raw documents
Distribution of the number of words after lemmatization and stop words removal Refer to caption
(a) Preprocessing for TF-IDF and LSTM
Distribution of the number of tokens after tokenization for Longformer Refer to caption
(a) Preprocessing for Longformer
Figure 6: Distribution of length of documents in Dataset #1, for different preprocessings.
3.2 Models for Case Classification
@@
In this section, we delve into the methodologies employed for analyzing our datasets, focusing on four distinct approaches: TF-IDF, LSTM, Longformer, and regex. Together, they faithfully represent some of the most common NLP methods.
@@
Term Frequency-Inverse Document Frequency (TF-IDF) is a vector embedding technique divided into two parts. In the “term-frequency” part, each document $d$ is embedded in the space $\mathbb{R}^{n}$, where $n$ is the number of different words on the corpora, and where the $i$-th component of the vector is the frequency that the $i$-th word of the corpora appears in the document. We denote it by $\text{TF}(i,d)$. The “inverse document frequency” part is another vector of $\mathbb{R}^{n}$, with $i$-th coordinate given by the logarithm
@@
By itself, TF-IDF is only an embedding, so it must be coupled with some classification algorithm for us to perform Case Classification. In this article, we will use three kinds of classification models for the embedded documents: linear Support Vector Machine (SVM), logistic regression, and random forests. We will not explicitly describe these classifiers here, but refer the interested reader to [28].
In practice, we trained the models on a binary task on Dataset #1 (described in Section 3.3), where 90% of the data was used for training and 10% for testing. Based on a validation set, the models were trained to maximize the $F_{1}$ score, from grid search on specific ranges of parameters ($\{1,5,10,50\}$ for SVM, $\{10,10^{2},10^{3},10^{4}\}$ for logistic, and $\{10,10^{2},10^{3}\}$ trees in the forest), through cross-validation with 5 folds.
@@
A recurrent network architecture was also implemented for this task, given the well-known capacity of these models to work with textual data, especially for longer texts [28, 51]. The implemented neural networks were composed of three layers: one embedding layer, followed by two layers consisting of 128 LSTM (Long Short-Term Memory) neurons each, all using relu activation functions. For the embedding layer, tensorflow was used, where the vocabulary (i.e., set of unique words in the corpora) was restricted to the 10,000 most common terms, using 1,000 out-of-bucket embedding for the rest of them. Moreover, because the architecture expects a constant input size of documents, only the first 4096 tokens of each document were considered, using additional masking for shorter documents. Although naive, this approach is common when dealing with long text analysis in neural networks. Finally, the output consisted of a single neuron followed by a sigmoid non-linearity. We stress that no pre-training of the network’s weights was used.
[…]
Longformer
It is a transformer-based architecture designed to efficiently handle long sequences by incorporating a modified self-attention mechanism that scales linearly with sequence length [4]. For sequence classification tasks, Longformer uses a global attention token that attends to all parts of the sequence, while each token within the sequence only attends to a local window around itself. This approach effectively reduces the computational overhead compared to standard full-attention transformers [69], enabling the model to process longer documents without sacrificing critical contextual information. As with other transformer models, Longformer includes layers of multi-head self-attention, feed-forward networks, and layer normalization.
In this article, we utilized the publicly available longformer-base-4096 model released by Allen Institute for AI (allenai) [^5] Pretrained model available at https://huggingface.co/allenai/longformer-base-4096. This model was pre-trained on a large corpus of English texts and fine-tuned in our experiments for sequence classification on Dataset #1. Even though its pre-training data is in English, and our data in Portuguese, the underlying learned representations of language structure, context, and semantic relationships are expected to transfer well to other languages.
More specifically, for each BP, the model was fine-tuned for 5 epochs (after grid search with 1, 3, and 5 epochs) through weighted binary cross-entropy loss (as LSTM), with a learning rate of $5\times 10^{-5}$ (the best in the range $\{1,5\}\times\{10^{-3},10^{-4},10^{-5}\}$), and a batch size of 64 (chosen among powers of two between 2 and 512) via gradient accumulation. The token limit we chose is 4096, the maximum for longformer-base-4096.
@@
Regular expression (regex) stands out as a commonly used tool among professionals for retrieving similar cases. In particular, it constitutes the search engine of the official STF website. We designed a regex search for the topics encompassed by the five BPs by selecting, with the help of legal expertise, the most important words in their statements (refer to Sections 5.2, 5.3, 5.4, 5.5 and 5.6 for the statement and juridical context of each BP, or to Section 2.1 for an overview). We draw the reader’s attention to the fact that, as our models are intended to detect, in a subsequent section, the documents before the creation of the BPs, we do not use the words Súmula Vinculante or any specific mention of a law. More precisely, we consider:
@@
We emphasize that these regular expression searches are fairly elementary, and could potentially be improved by Named-entity recognition (NER) techniques. However, they are quick to perform, and highly explicable, allowing direct interpretation of the results. In addition, as we will see during the validation, they achieved surprisingly correct scores (see Table 2).
Running times
Training the TF-IDF models and applying them to Datasets #1 and #2 did not require significant computational resources and were performed on a CPU. In contrast, the LSTM and Longformer models required a GPU for both training and prediction. Among the models, Longformer was the most resource-intensive. For each BP, the training process took approximately 14 hours [^6] On a system equipped with an Intel(R) Xeon(R) Gold 6230 CPU, 192GB of RAM, and an Nvidia(R) Dual Quadro RTX6000 GPU, using the parameters detailed in the corresponding section above. The primary bottleneck in training stemmed from the document lengths (4096 tokens). On the other hand, prediction on Dataset #2 (containing 634,068 documents), if carried out without tricks, would take around 31 hours (approximately 175 milliseconds per document), using the maximum batch size supported by the GPU, which is 16. To reduce this time, we employed two optimization strategies. First, we utilized the fast vectorization capability natively included with longformer-base-4096 and enabled FP16 computational precision, which reduced the processing time by 55% (i.e., 78 milliseconds per document). Second, we implemented dynamic batching: instead of grouping documents into fixed-size batches of 64 and padding (or truncating) them to 4096 tokens, we ordered documents by length and dynamically created the largest possible batches. More precisely, we ensured that, after padding to the length of the largest document in the batch, the total token count did not exceed $16\times 4096$. This further reduced the processing time by 33% (or 53 milliseconds per document). With these optimizations, the prediction of Dataset #2 using a Longformer model was completed in approximately 9 hours.
@@ 3.3 Validation
Our first analysis focuses on Dataset #1, a collection of documents labeled by the binding precedent they cite. We are, therefore, dealing with a binary classification task. As we will see in this section, and as is expected from the NLP literature, all the models perform well on this task. That said, we remind the reader that the main objective of this article lies in the analysis of Dataset #2 (in Sections 4 and 5), which is unlabeled and presents a more challenging scenario with different outcomes.
@@
Some preprocessing was applied to the documents before training. For both the TF-IDF and LSTM models, this encompassed the lemmatization of the texts, the removal of Portuguese-specific accents, and of usual stop words (e.g., a, o, um, mas). Because Longformer does not expect this sort of preprocessing, raw texts were given.
Additionally, we want to avoid the models from learning textual cues that mark direct citations to BPs, such as the very term Súmula Vinculante or similar. Therefore, regular expressions were used to substitute these terms with empty strings. Additionally, as our interest consists in detecting possible uses of BPs before their publications, dates were removed. Once more, Longformer models receive a slightly different masking process than the others: each BP citation is identified using regular expressions and is replaced by <mask> tokens, keeping the same length as the original text.
@@
A word of caution is warranted regarding how the data are split into training and test sets. As opposed to a multiclass dataset, where each document would be associated with one label, our dataset is multilabeled, in the sense that certain documents can be associated with several labels. To split such a dataset, it would be ideal to reproduce the same distribution of combinations of labels in the training and the test set. To achieve this, we use the iterative stratification algorithm [^7] Multilabel data stratification is implemented in skmultilearn http://scikit.ml/stratification.html presented in [55, 62].
Classification task
From a methodological point of view, we emphasize that, in this section, our approach consists of training models for the recognition of citations to the binding precedents. Furthermore, the models are trained on Dataset #1, where explicit references to BPs have been removed; this will ensure that, when later applied to Dataset #2, the models will not simply be able to detect documents that cite one of the five BPs, but more generally documents that, from a legal point of view, fall within the scope of the BP.
In order to detect documents citing one of the five BPs considered, we will train five independent models, one for each BP, addressing the binary task related to the BP. In other words, we have transformed the multiclass classification task into five binary classification tasks. However, another approach, common in such problems, is worth mentioning: training only one multiclass classifier, capable of outputting the BP detected among the five, or else a null class if no citation is detected. That said, this latter point of view is not well suited to our problem, for two reasons. Firstly, our problem is inherently multilabel: as we saw in Table 1, 4% of the documents in Dataset #1 cite two BPs or more. By training separate binary classifiers, we enable the possibility of potentially assigning several BPs to a single document.
Secondly, and perhaps more importantly, we do not want the detections of one BP to influence another. Indeed, although the BPs deal with distinct legal themes, the limits of their scope in practice can be blurred, sometimes overlapping with one another. A unified multilabel model risks narrowing their scope, which is not the case with isolated binary classifiers. This phenomenon will be observed, for example, in the analyses of BPs 17 and 37 (in Sections 5.4 and 5.6). While one addresses late payment penalties and the other deals with salary adjustments, there are cases that intertwine these two themes, or that rely on common legal arguments (such as isonomy).
@@
$\mathrm{Recall}$ $=\mathrm{TP}/\mathrm{P},$
@@
Intuitively, accuracy measures the ratio of correct predictions, while precision measures the accuracy of pointing to class 1, and recall measures how much of class 1 is being correctly predicted. The $F_{1}$ score is the harmonic mean of recall and precision. We point out that because our classes are highly imbalanced (with many more documents in class 0 than in class 1, as seen in Table 1), the accuracy is of limited value. Our reference is the $F_{1}$ score, which specifically measures how well class 1 is predicted.
Although our models ultimately produce a binary decision, some internally generate a continuous output, that can be interpreted as a probability of belonging to class 1. For instance, LSTM’s native output is a value $p$ between 0 and 1, and one classifies a document as class 1 if $p\geq 0.5$. In contrast, Longformer’s raw output is a logit, represented as a pair $(p,q)$ that can be greater than 1 or less than 0. One typically applies a softmax activation to this logit, producing a new pair $(p^{\prime},q^{\prime})$ in the interval $[0,1]$. The second value, $q^{\prime}$, is then interpreted as the probability of belonging to class 1, and one predicts the class by checking if $q^{\prime}\geq 0.5$. For the classifiers paired with TF-IDF, only logistic regression natively uses a probability value, for it is inherently a probabilistic model. Each document results in a value $p$ and is classified as class 1 if $p\geq 0.5$. We can thus directly request this probability output from the model. By contrast, SVM and random forest do not have a native notion of probability, though one can force them to return a continuous output. In the case of SVM, a natural continuous quantity is the signed distance $d$ from the margin; a document is said to belong to class 1 if $d\geq 0$. One can transform $d$ into a probability via Platt scaling, obtaining a new value $d^{\prime}$, and then classify according to $d^{\prime}\geq 0.5$. Note that the conditions $d\geq 0$ (“pure” SVM prediction) and $d^{\prime}\geq 0.5$ (“calibrated probability”) are not equivalent: they may disagree slightly on near-boundary samples. However, the results are generally very similar. Last, random forest can also straightforwardly provide a probability: instead of returning the raw majority-vote class label from the ensemble of decision trees, it averages the fraction of training samples in each leaf that belongs to the positive class. The final decision is again made by checking if $p\geq 0.5$.
Based on these probability outputs, one could fine-tune the thresholds — rather than always using 0.5 — to potentially achieve better predictions. The Area Under the Precision-Recall Curve (AUPRC) swaps a classification threshold between extreme values, measuring precision and recall for each threshold, and creates a precision vs. recall curve. The area under this curve establishes a common performance metric.
The results of all models are presented in Table 2. We draw the reader’s attention to the fact that regex has an undefined AUPRC score since there is no natural way of associating its predictions with probabilities. Once again, we emphasize that the main goal of this article is to analyze binding precedent efficiency through case classification on Dataset #2. This entire section, devoted to comparing the models on Dataset #1 (our reference dataset for validation, since it is labeled), is only one step in our overall approach. In Section 4, we will examine how the models perform on Dataset #2, where the outcomes may not align perfectly with the scores reported below.
@@
11 TF-IDF+SVM 98.2 98.8 97.6 98.8
TF-IDF+logistic 96.9 100 94.0 98.8
TF-IDF+forest 91.5 100 84.3 98.6
LSTM 91.1 89.5 92.8 94.1
Longformer 95.1 97.5 92.8 96.2
regex 91.6 98.6 85.5 $\times$
14 TF-IDF+SVM 99.1 100 98.1 99.9
TF-IDF+logistic 98.7 100 97.5 99.9
TF-IDF+forest 96.2 98.7 93.8 99.8
LSTM 94.9 91.8 98.1 99.6
Longformer 98.1 98.7 97.5 99.4
regex 78.7 98.1 65.6 $\times$
17 TF-IDF+SVM 98.8 100 97.7 100
TF-IDF+logistic 98.8 100 97.7 100
TF-IDF+forest 95.8 100 91.9 100
LSTM 95.2 98.8 91.9 97.9
Longformer 98.8 100 97.7 98.6
regex 92.9 94.0 91.9 $\times$
26 TF-IDF+SVM 98.9 100 97.9 99.1
TF-IDF+logistic 97.3 100 94.7 98.9
TF-IDF+forest 85.5 100 74.7 98.9
LSTM 97.8 100 95.8 98.7
Longformer 94.6 97.8 91.6 97.0
regex 76.9 98.4 63.2 $\times$
37 TF-IDF+SVM 96.9 100 94.0 99.6
TF-IDF+logistic 96.1 100 92.5 99.6
TF-IDF+forest 91.4 100 84.2 99.5
LSTM 96.5 99.5 93.7 98.2
Longformer 97.2 99.2 95.2 99.9
regex 86.7 93.3 80.9 $\times$
Table 2: Test scores for the task of binary classification in Dataset #1. For each BP and each model, we report the metrics $F_{1}$, precision, recall, and AUPRC (in percentage). The binary scores ($F_{1}$, precision and recall) are computed from the models’ raw outputs, whereas the AUPRC is computed from their probability outputs. As described in the text, the models’ binary outputs are obtained by applying a threshold of 0.5 to their probability outputs. For each BP, the best score in each column is shown in bold.